OpenStudy (anonymous):

How do I determine the value of x that makes x-2, x+2, 5, 9,... (arithmetic sequence)?

OpenStudy (hoblos):

9-5 = 4 =>5 -(x+2) =4 =>x=-1

OpenStudy (anonymous):

common difference is 4 x+2 -(x-2) = 4 5-(x+2) = 4 5-x-2 =4 3-x = 4 x = -1